OpenStudy (anonymous):

I need help with this please! -3=(-6+x)/2 -5=(-3+y)/2 show step by step how to solve this please

jimthompson5910 (jim_thompson5910):

I'll solve the first equation and let you solve the second equation -3=(-6+x)/2 -3*2 = -6+x ... Multiply both sides by 2. -6 = -6 + x -6 + 6 = -6 + x + 6 ... Add 6 to both sides. 0 = 0 + x 0 = x x = 0 So the solution to -3=(-6+x)/2 is x = 0

OpenStudy (anonymous):

okay so basically -10=-3+y -13=y am i correct?

jimthompson5910 (jim_thompson5910):

-10 = -3 + y is correct

jimthompson5910 (jim_thompson5910):

but when you solve -10 = -3 + y for y, you won't get y = -13

jimthompson5910 (jim_thompson5910):

-10 = -3 + y -3 + y = -10 y = -10 + 3 y = ???

OpenStudy (anonymous):

oh yeah forgot. -7

jimthompson5910 (jim_thompson5910):

correct, y = -7

OpenStudy (anonymous):

okay but when u said multiply by both sides, u only multiplied one side?

jimthompson5910 (jim_thompson5910):

when you multiply (-6+x)/2 by 2, the 2 in the denominator will cancel

jimthompson5910 (jim_thompson5910):

start with this 

jimthompson5910 (jim_thompson5910):

then multiply it by 2 to get